### v3

# Conditional Video Generation for High-Efficiency Video Compression

Abstract
Perceptual studies demonstrate that conditional diffusion models excel at reconstructing video content aligned with human visual perception. Building on this insight, we propose a video compression framework that leverages conditional diffusion models for perceptually optimized reconstruction. Specifically, we reframe video compression as a conditional generation task, where a generative model synthesizes video from sparse, yet informative signals. Our approach introduces three key modules: (1) Multi-granular conditioning that captures both static scene structure and dynamic spatio-temporal cues; (2) Compact representations designed for efficient transmission without sacrificing semantic richness; (3) Multi-condition training with modality dropout and role-aware embeddings, which prevent over-reliance on any single modality and enhance robustness. Extensive experiments show that our method significantly outperforms both traditional and neural codecs on perceptual quality metrics such as Fréchet Video Distance (FVD) and LPIPS, especially under high compression ratios.
Introduction
The exponential growth of video content across streaming platforms, social networks, teleconferencing, and augmented reality applications has created unprecedented demand for effective compression techniques. Current video compression standards, including H.266/VVC (Zhang et al., 2020) and AV1(Chen et al., 2020), have achieved substantial improvements through decades of engineering refinement, employing hybrid coding strategies that combine motion estimation, transform coding, and entropy modeling. However, these approaches rely on largely handcrafted components within rigid codec architectures, limiting their adaptability to diverse application requirements.
Most traditional and neural compression pipelines operate under a fundamental assumption: the pursuit of pixel-level fidelity to ensure reconstructed frames match the original input as closely as possible. While this approach suits applications requiring exact reproduction—such as scientific imaging or professional video editing—we argue that strict fidelity is often unnecessary for perceptual consumption scenarios. In applications like user-generated content, entertainment streaming, or virtual conferencing, perceptual consistency—visual coherence aligned with human perception—matters more than exact pixel reconstruction. Relaxing pixel-perfect accuracy requirements creates opportunities for aggressive compression while enabling new trade-offs between bitrate and perceptual quality. Data-driven approaches have emerged as promising alternatives to traditional codecs. Recent advances in neural image and video compression leverage encoder-decoder architectures and learned entropy models to achieve competitive rate-distortion performance. However, most methods remain constrained by deterministic reconstruction requirements and often exhibit suboptimal perceptual quality, particularly at low bitrates, manifesting as blurring, blocking artifacts, and color degradation.
Concurrently, generative models—especially diffusion models—have demonstrated state-of-the-art performance in image and video synthesis. This paradigm shifts focus from encoding pixel-level residuals to achieving content-faithful reconstruction under strict bitrate constraints by leveraging the strong priors of generative models and compact spatio-temporal guidance. While existing methods (Zhang et al., 2025; Wan, Zheng, and Fan, 2024; Wu et al., 2023) model spatio-temporal information via textual prompts, keyframes, or basic visual cues, we contend these representations are insufficient for high-fidelity video reconstruction. This limitation confines their applicability to narrow domains—such as human-face video (Chen et al., 2024), human-body video (Wang et al., 2023a, 2022), or small motion video (Yin et al., 2024) — where scene complexity remains constrained.
We explore this question through a diffusion-based compression framework that introduces three core innovations. First, we employ multi-granular signals that capture both static scene structure such as auto selected keyframes and semantic descriptions, and dynamic information including human motion, optical flow, and panoptic segmentation. Second, we design compact, transmission-efficient representations for these signals that serve as minimal yet perceptually informative inputs to the decoder. Third, we develop a multi-condition training strategy that incorporates signal dropout and role-aware embeddings, enabling the model to remain robust even when certain signals are unavailable or degraded.
We evaluate our method on standard video benchmarks using perceptual metrics including Fréchet Video Distance (FVD) and LPIPS. Results demonstrate substantial improvements over both traditional and learning-based codecs in perceptual quality, particularly at high compression ratios. These findings suggest that conditional generative models offer a promising paradigm for perception-centric video compression, where semantic compactness and visual plausibility supersede strict pixel accuracy.
To summarize, our main contributions are threefold:
  • First, we design an efficient and modular compression framework, incorporating multi-granular conditioning and compact, transmission-friendly representations to guide reconstruction with minimal overhead.
  • Second, we deploy a multi-condition training strategy enabling the model to remain robust even when certain signals are unavailable or degraded.
  • Third, we demonstrate state-of-the-art perceptual performance on standard benchmarks, significantly outperforming traditional and neural codecs under high compression ratios, as measured by FVD and LPIPS.
Related Work
Video Compression
Video compression aims to reduce redundant information in video sequences while preserving critical visual content, enabling efficient storage and transmission across applications such as streaming services, video conferencing, and surveillance. With the success of deep learning in image compression (Mishra, Singh, and Singh, 2022), neural video compression has emerged as a prominent research area, leveraging neural networks to optimize rate-distortion trade-offs. Residual coding-based approaches (Choi and Bajić, 2019) generate predicted frames from previously decoded ones and encode the residual between predicted and current frames. However, their reliance on simple subtraction for inter-frame redundancy reduction leads to suboptimal performance. Lu et al. (Li, Li, and Lu, 2021) pioneered this direction by replacing traditional codec modules with neural networks in an end-to-end framework. The DCVC series (Li, Li, and Lu, 2021) exemplifies this paradigm, with DCVC-DC (Li, Li, and Lu, 2023) and DCVC-FM (Li, Li, and Lu, 2024) outperforming traditional codecs like ECM (Karadimitriou, 1996). While these neural video compression methods have made significant strides in improving rate-distortion performance as measured by pixel-level metrics, they often overlook the perceptual quality of the reconstructed videos, which is crucial for human viewing experience. This gap between pixel-level metrics and perceptual quality led Blau et al. (Blau and Michaeli, 2019) to highlight a ”rate-distortion-perception” trade-off, spurring research into perceptual video compression. Methods in this domain optimize for visual quality by integrating perceptual losses (e.g., LPIPS (Zhang et al., 2018)) into rate-distortion objectives. GAN-based approaches further enhance realism: Mentzer et al. (Mentzer et al., 2022) framed the compressor as a generator trained adversarially to reconstruct detailed videos; Zhang et al. (Zhang et al., 2021) extended DVC (Lu et al., 2019) with a discriminator and hybrid loss for balanced rate, distortion, and perception.
Controllable Video Generation
Controllable video generation (as defined by VAST (Zhang et al., 2024)) aims to synthesize videos adhering precisely to external conditions (appearance, layout, motion) while maintaining spatiotemporal consistency. Early methods extended image diffusion models with text guidance, enabling creative generation but lacking fine-grained detail and motion control. Subsequent approaches incorporated stronger conditions: Image animation methods used initial frames but often produced static results. Methods using low-level dense signals (e.g., depth or edge sequences: Gen-1 (Esser et al., 2023), ControlVideo (Zhao et al., 2025), VideoComposer (Wang et al., 2023b)) improved control but proved impractical. Object trajectory or layout control emerged via strokes (DragNUWA (Yin et al., 2023)), coordinates (MotionCtrl (Wang et al., 2024b)), or bounding boxes. Training-based trajectory methods (TrackGo (Zhou et al., 2025)) were costly with limited gains, while training-free attention manipulation (FreeTraj (Qiu et al., 2024)) suffered inaccuracies. Critically, existing methods remain fragmented: each control modality typically requires specialized inputs or architectural changes, highlighting the need for unified, adaptable frameworks.
Diffusion Models for Video Compression
Diffusion-based compression has advanced rapidly in the image domain, laying groundwork for video applications. Wu et al. (Wu et al., 2023) transmitted sketches and text descriptions to guide diffusion-based reconstruction, while Careil et al. (Careil et al., 2023) used vector-quantized latents and captions for decoding. Relic et al. (Relic et al., 2025) optimized efficiency by framing quantization noise removal as a denoising task with adaptive steps. However, extending these advances to video faces key challenges: integrating foundational diffusion models into existing video coding paradigms (e.g., conditional coding) without disrupting efficiency, mitigating slow inference, and enabling multi-bitrate support for varying latent distortion levels. To address these, this work focuses on diffusion-based video compression, emphasizing dynamic and static condition controls to balance compression ratio and perceptual quality. Dynamic controls leverage temporal dependencies (e.g., motion vectors, previous frames) to capture video dynamics, while static controls incorporate semantic cues (e.g., scene categories, texture priors). By harmonizing these conditions, the proposed method enhances reconstruction fidelity across bitrates while preserving coding efficiency—bridging the gap between diffusion-based image compression success and unmet needs in video coding.
Figure 1: Our framework processes input videos through three sequential stages: First, a keyframe selection module partitions the video into consecutive clips. Second, clip-specific conditions are compressed—the first frame, last frame, and textual descriptions via entropy coding while segmentation sequences, human motion data, and optical flow are converted to compact representations. Finally, at the decoder, a controllable diffusion model reconstructs each clip using all decompressed conditions to generate the output video.
Method
Fig. 1 illustrates the framework of our proposed method. Our video compression approach comprises three main stages:
Keyframe Selection & Clip Segmentation. The original video is processed by a keyframe selection module. This partitions the video into consecutive clips. Each clip is defined such that its first and last frames are designated as keyframes, and serves as an independent unit for transmission.
Conditional Feature Extraction & Compression. For the intermediate frames within each clip, we extract conditional representations. These include textual descriptions, segmentation maps, human motion sequences, and optical flow sequences. These representations are subsequently converted into a compact form.
Conditional Frame Generation at Decoder. For each clip, both the compressed keyframes and the compact conditional representations are then transmitted over the network to the decoder. At the decoder, the intermediate frames of each clip are reconstructed using a pre-trained multi-conditional diffusion model.
1. Key Frame selection
Our keyframe selection strategy employs dual-criterion detection to balance compression efficiency and reconstruction quality. Formally, given video sequence $\mathcal{V}=\{f_{1},f_{2},\dots,f_{T}\}$ with $T$ frames, we identify keyframes through:
  1. (i)
    Shot boundary detection: Frame $f_{i}$ is selected when identified as a shot transition frame using TransNetV2 (Souček and Lokoč, 2020). We compute shot transition probability $p_{i}=\mathcal{T}(f_{i})$ where $\mathcal{T}$ denotes the pretrained TransNetV2 model, and designate $f_{i}$ as a keyframe when $p_{i}>0.5$.
  2. (ii)
    Fixed-interval sampling: Frame $f_{i}$ is selected if $i-i_{\text{prev}}\geq w$, where $w$ is a tunable hyperparameter and $i_{\text{prev}}$ denotes the previous keyframe index. Smaller $w$ values increase keyframe density (improving reconstruction quality at higher bitrates) while larger $w$ reduces transmission overhead.
Algorithm 1 Keyframe-based Clip Segmentation
1:Video frames $\mathcal{V}=\{f_{1},f_{2},\dots,f_{T}\}$,
2:   Hyperparameter $w$,
3:   Shot detector $\mathcal{T}$
4:Clips $\mathcal{C}=\{[start_{m},end_{m}]\}_{m=1}^{M}$
[…]
8:for $i\leftarrow 1$ to $T$ do
9:  if $(i-\mathit{last\_key})\geq w$ or $\mathcal{T}(f_{i})>0.5$ then
10:   $\mathcal{K}\leftarrow\mathcal{K}\cup\{i\}$
11:   if $(i-\mathit{last\_key})\geq w$ then
12:     $\mathit{prev\_type}\leftarrow\texttt{interval}$
13:   else
14:     $\mathit{prev\_type}\leftarrow\texttt{shot}$
15:   end if
16:   $\mathit{last\_key}\leftarrow i$
17:  end if
18:end for
[…]
22:for $j\leftarrow 1$ to $|\mathit{keys}|-1$ do
23:  $s\leftarrow\mathit{keys}[j-1]$
24:  $e\leftarrow\mathit{keys}[j]$
25:  if $\mathit{prev\_type}=\texttt{shot}$ then
26:   $s\leftarrow s+1$ $\triangleright$ Offset for shot boundary
27:  end if
28:  $\mathcal{C}\leftarrow\mathcal{C}\cup\{[s,e]\}$
29:  $\mathit{prev\_type}\leftarrow\text{type of }\mathit{keys}[j]$
30:end for
The combined keyframe indices $\mathcal{K}=\{k_{0},k_{1},\dots,k_{M}\}$ partition $\mathcal{V}$ into $M$ clips $\{\mathcal{C}_{1},\mathcal{C}_{2},\dots,\mathcal{C}_{M}\}$. Crucially, clips originating from shot boundaries offset segment boundaries to avoid assigning transition frames to adjacent clips. This prevents generation artifacts during scene transitions where inter-frame discontinuities exceed generative model capabilities. The clip segmentation procedure is formulated as Algorithm 1.
2. Conditional Feature Extraction
Inspired by the storyboard paradigm in VAST (Zhang et al., 2024), spatiotemporal control signals significantly enhance motion dynamics, spatial alignment, and temporal consistency - critical factors for perceptual fidelity in video reconstruction. Besides the textual descriptions obtained through the Starlight Multimodal Understanding Large Model (TODO), we select three complementary conditional representations whose compact forms balance reconstruction quality against transmission bandwidth, adhering to the VAST(Zhang et al., 2024) principle that increased conditional complexity improves reconstruction fidelity.
Figure 2: Our diffusion model converts optical flow, segmentation, and motion into visual modalities, encodes them via VAE, and concatenates the latent codes with noise as input to the diffusion backbone.
Segmentation Sequences
Segmentation sequences provide crucial geometric scaffolding by preserving object boundaries and spatial relationships across frames. This explicit structural prior prevents shape distortion during generation and maintains consistent object interactions throughout temporal transitions. We extract per-frame panoptic segmentation map using Mask2Former (Cheng et al., 2021). To achieve high compression, we first extract its external contour using border tracing, and then Approximate each contour with $n$-th order Bézier curves:
$B(t)=\sum_{i=0}^{n}\binom{n}{i}(1-t)^{n-i}t^{i}P_{i},\quad t\in[0,1]$
where $P_{i}$ are optimized control points. Finally, we retain only the $N$ longest contours per frame. Fig. 3(a) shows the original segmentation map and the fitting results obtained with different values of $N$, where 8th-order Bézier curve is applied. This compact representation preserves essential object topology while largely reducing storage since only the Bézier parameters need to be saved.
Human Motion Representation
Human motion sequences are essential for human-centeric videos(Zhang et al., 2024) since it significantly reduces artifacts in articulated movements and maintains natural temporal coherence during complex actions like walking or dancing. We represent human motion using 3D SMPL sequences (Loper et al., 2015). Then, we select 21 joints that represents human kinetics and project those 3D joints to 2D image-coordinates. By calculating the proportion of the area occupied by 2D joints in the image coordinates, we set different thresholds to filter out small human poses. Fig. 3(b) shows the human motion results obtained with different area-threshold $\xi$}. The 2D joints coordinates are transmitted over the network.
Optical Flow Representation
Optical flow fields explicitly encode dense displacement vectors between pixels, providing critical motion guidance. We compute optical flow using RAFT (Teed and Deng, 2020). However, the cost of transmitting pixel-level optical flow is prohibitively expensive. We believe that the representation of optical flow should be based on blocks of pixels rather than individual pixels. Specifically, we define sampling stride $l$ (hyperparameter controlling compression ratio) and then sample flow vectors at regular intervals.
$\mathcal{G}=\left\{\mathcal{G}(x,y)\mid x=\lfloor i\cdot l\rfloor,y=\lfloor j\cdot l\rfloor\right\}$
where $i=1,2,\dots,\lfloor H/l\rfloor$, $j=1,2,\dots,\lfloor W/l\rfloor$, $\mathcal{G}$ is the extracted optical flow map.
Under large sampling strides, bilinear interpolation produces significant errors in recovered optical flow fields. Therefore, we employ a flow-arrow visualization approach where arrow direction indicates motion orientation at sampled points and arrow length represents flow magnitude. Fig. 3(c) demonstrates this flow visualization across different sampling strides.
Compression Calculation
For a clip, the first and last frames are encoded into a bitstream using a state-of-the-art image compression method, such as LIC (Li et al., 2025), along with text, with a size set to QQKB. The representations of the remaining three conditions are directly encoded in numeric form, with each number represented using bfloat16 (2 Bytes). The compressed bitrate(KBps) of our framework is calculated per video clip as:
$\begin{split}R=\frac{Q\cdot\text{fps}}{T}&+\frac{2\cdot\text{fps}}{1024}\cdot\Bigg{[}\\ &\quad\phi(\xi)\cdot 21\cdot 2+2\left\lfloor\frac{H}{l}\right\rfloor\left\lfloor\frac{W}{l}\right\rfloor+2N(n+1)\Bigg{]}\end{split}$ (1)
  • $Q$: Size of compressed first frame and last frame + text (KB)
  • $T$: Clip length
  • fps: Frame rate (frames/sec) of the clip
  • $k$: The number of people pose remaining under threshold $\phi({\xi})$
  • $H,W$: Frame dimensions (pixels)
  • $l$: Flow sampling stride
  • NN: The number of Bézier curves
  • $n$: Order of Bézier curve (default: 8)
Figure 3: Visualization of optical flow, human motion, and segmentation representations alongside their compact forms at varying bitrate thresholds.
3. Conditional Frame Generation at Decoder
Our controllable diffusion model is built upon the pretrained FL2V (first&lastframe-to-video) diffusion model VAST (Zhang et al., 2024). Firstly, the condition signals, i.e. segmentation/human motion/optical flow are convert to dense visual modalities $V=\{V_{seg},V_{motion},V_{flow}\}\in R^{T*H*W*3}$, as shown in Fig. 3. Given the paried visual modalities, we first encode them into a latent space using a pretrained 3D causal VAE encoder $\epsilon$ (Zhang et al., 2024).
$x_{m}=\epsilon(V_{m}),m\in\{flow,motion,seg\}$
Next, we concatenate these latents along the channel dimension with the latent noise, forming the input of the diffusion transformer. During training, a key challenge arises from modality entanglement: the model often over-relies on dominant conditions (e.g., segmentation) while neglecting others. To enforce balanced utilization, We apply random dropout for each condition with a dropout ratio of 0.3. For any condition subjected to dropout, its visual representation is replaced with a zero-valued (all-black) image sequence.
However, random dropout introduces an another problem: role ambiguity. The model cannot distinguish between intentionally absent conditions and dropout-induced zeros. To address this, we introduce an adaptive control strategy that dynamically assigns roles to different modalities. We introduce a modality embedding $e_{m}$ that differentiates between dropout ($e_{d}$) and conditioning ($e_{c}$) roles, which can be directly added to the diffusion model input.
$\mathbf{e}_{m}=\begin{cases}\mathbf{e}_{d},&\text{if m is subjected to dropout}\\ \mathbf{e}_{c},&\text{conditioning}\end{cases}$ (2)
This strategy enables flexible and efficient control, allowing the model to seamlessly adapt to different tasks without requiring separate architectures for each modality. The model is capable of learning a joint representation of multiple conditions while simultaneously avoiding over-reliance on certain conditions.
Experiments
To evaluate the performance of our method and state-of-the-art (SOTA) methods on the video compression task, we conduct a series of experiments and empirical studies on our collected data, which contains both open-source datasets like Koala-36M (Wang et al., 2024a) and the data we crawled from the Internet.
Experimental Settings
Datasets. We group our collected data into four major categories: dancing, gymnastics, diving, and scenery videos. For scenery videos, we use optical flow and segmentation as conditional features, and set the motion representation as a zero-valued (all-black) image sequence. For other human-centric videos, we additionally include motion data in the conditions.
Category # Raw Videos # Avg. Duration (m)
Dancing 102,235 1.8
Gymnastics 49,925 15.8
Diving 56,934 20.2
Scenery 235,622 8.9
Table 1: Statistics of the preprocessed test set.
Data Preparations. For data preparations, we adopt a filtering strategy that filters out videos that are labeled ”poor” (i.e., low resolution, low aesthetic score, etc.) to reduce noise and ensure the quality of the datasets. We train and tune all models on the internal data five times, choose the best epoch on the validation set for each training as the model to be tested on the test set, and report the average experimental results. Table 1 summarizes detailed statistics of our test set after pre-processing.
Evaluation Metrics. We use two commonly used metrics for evaluation: Fréchet Video Distance (FVD) and Learned Perceptual Image Patch Similarity (LPIPS), as they better align with human perception compared to traditional measures such as Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index Measure (SSIM).
Given the distributions of real and generated videos $P$ and $Q$, we have
$\text{FVD}(P,Q)=\left\|\bm{\mu}_{P}-\bm{\mu}_{Q}\right\|_{2}^{2}+\text{Tr}\left(\bm{\Sigma}_{P}+\bm{\Sigma}_{Q}-2\sqrt{\bm{\Sigma}_{P}\bm{\Sigma}_{Q}}\right),$ (3)
where $\bm{\mu}_{P},\bm{\mu}_{Q}$ are the mean vectors of real videos and generated videos in the feature space of a pre-trained network (e.g., I3D), $\bm{\Sigma}_{P},\bm{\Sigma}_{Q}$ are the corresponding covariance matrices, $\text{Tr}(\cdot)$ denotes the trace of a matrix, $\sqrt{\bm{\Sigma}_{P}\bm{\Sigma}_{Q}}$ is the matrix square root of the product of the covariance matrices (computed via eigenvalue decomposition).
$\text{LPIPS}(x,x_{0})=\sum_{l}\frac{1}{H_{l}W_{l}}\sum_{h,w}\left\|w_{l}\odot\left(\hat{y}_{hw}^{l}-\hat{y}_{0hw}^{l}\right)\right\|_{2}^{2},$ (4)
where $l$ is the hidden features of the $l$-th layer in the pre-trained networks (like VGG and AlexNet), $\hat{y}^{l}$ and $\hat{y}_{0}^{l}$ are the normalized activation values, $w_{l}$ is the weight of the $l$-th layer features (obtained from human perception studies).
Baselines. To demonstrate the effectiveness of the framework, we compare the proposed framework with both traditional video compression standards, H.264 and H.265, and SOTA video compression methods:
  • H.264: Also known as Advanced Video Coding (AVC), it is a widely adopted video compression standard that significantly improves efficiency while maintaining good video quality, making it ideal for bandwidth-constrained applications.
  • H.265: Also known as High Efficiency Video Coding (HEVC), it is the successor to H.264, designed to further reduce file sizes by up to 50% at the same quality level, making it especially beneficial for 4K, 8K, and HDR content, where bandwidth and storage savings are critical.
  • DCVC-RT  (Jia et al., 2025): It is the latest flagship model of the DCVC family, achieves real-time neural video coding by prioritizing operational efficiency—not just computation—to achieve unprecedented speed without sacrificing compression. Unlike existing codecs, DCVC-RT eliminates bottlenecks like memory I/O and excessive function calls through implicit temporal modeling and single low-resolution latents.
All these baselines can be divided into two categories: (1) traditional video compression methods (H.264, H.265); and (2) neural video compression methods (DCVC-RT).
Compression Settings. To demonstrate the effectiveness of our framework comprehensively, we design multiple compression settings, as shown in Table 2. Specifically, for level 0 setting, we only use the first and the last frame of the clip, with no additional conditions provided.
Level Segmentation Human Motion Optical Flow
Level 0 N/A N/A N/A
Level 1 $N=10$ $\xi=1/5$ $l=128$
Level 2 $N=20$ $\xi=1/8$ $l=96$
Level 3 $N=30$ $\xi=1/10$ $l=64$
Table 2: Different compression settings.
Hyperparameter Settings.
For a fair comparison, all methods are implemented with Pytorch 2.4 in Python 3.9.13 and learned with Adam optimizer (Kingma and Ba, 2017). We conduct our experiments on a single Linux server with 2 Intel(R) Xeon(R) CPU Platinum 8558 @2.10 GHz, 2 TB RAM, and 8 NVIDIA H100 (80 GB of graphic memory each). We fine-tune the pre-trained VAST-10B for 1 epoch with a learning rate of $2\times 10^{-5}$ and a batch size of 8. We tuned the parameters of all methods over the validation set.
Overall Performance
Table 3 and Fig. 4 summarize the overall performance of all models. Here, we make the following observations.
Method Level 0 (bpp = 0.0024) Level 1 (bpp = 0.0066) Level 2 (bpp = 0.0099) Level 3 (bpp = 0.0183)
FVD (↓) LPIPS (↓) FVD (↓) LPIPS (↓) FVD (↓) LPIPS (↓) FVD (↓) LPIPS (↓)
H.264 3835 0.5009 3085 0.4877 1751 0.4014 1258 0.3517
H.265 3467 0.4701 2344 0.4546 1447 0.3767 923 0.2732
DCVC-RT 3533 0.4889 2305 0.4479 1482 0.3502 641 0.2075
Ours 2415 0.4522 2283 0.4240 1347 0.3488 803 0.2566
Table 3: The overall performance of different methods on the test set. Here, we report the FVD and LPIPS among actual videos and different methods at different compression ratios. The best results are highlighted in boldface.
First, our diffusion-based video compression framework (conditioned on human motion, canny edges, and optical flow) outperforms traditional codecs (H.264/H.265) and neural compression baselines (e.g., DCVC-RT) in nearly all bitrate settings. Quantitative metrics (FVD and LPIPS) show improvements of 15–30%, particularly in extreme low bitrate settings, confirming that the generated videos better preserve perceptual quality , avoiding blocking artifacts (common in traditional codecs) and over-smoothed textures (typical of neural methods).
Second, while higher compression ratios lead to lower objective scores, the generative nature of diffusion ensures graceful degradation in perceptual quality. Even at level 1 compression where bpp is less than 0.007, key motion and semantics remain recognizable (see visualized results), making the method viable for bandwidth-constrained applications (e.g., mobile streaming, surveillance).
Last, current decoding is slower (~2 fps) than traditional codecs, but optimizations (latent-space compression, 1/8-resolution flow maps) should improve speed. Future work on distillation and hardware acceleration could enable real-time deployment.
Figure 4: Rate-distortion (perception) performance on the test set.
Ablation Study
We conduct a comprehensive ablation study to evaluate the contribution of each condition in our framework: segmentation sequences (Seg), human motion (Motion), and optical flow (Flow). Table 4 compares the performance of different condition settings across two compression levels. Here, we can make the following observations:
Table 4: The ablation study of condition designs. The best results are highlighted in boldface.
Condition Level 1 Level 3
FVD (↓) LPIPS (↓) FVD (↓) LPIPS (↓)
w/o Seg 2506 0.4384 942 0.2924
w/o Motion 2724 0.4510 1012 0.3075
w/o Flow 2352 0.4280 863 0.2724
full model 2283 0.4240 803 0.2566
The ablation study (Table 4) reveals how conditioning signals interact with bitrate constraints.
First, across both compression levels, the absence of human motion (w/o Motion, row 2) causes the largest performance drop at low bitrates compared with the full model (row 2 vs. row 4). For example, at Level 1, removing Motion alone increases FVD and LPIPS significantly, suggesting its critical role in preserving temporal coherence. This aligns with our test set’s human-centric content (e.g., interviews, sports), where body movements dominate perceptual quality. Notably, motion’s impact remains important but less dominant at Level 3, as increased bitrates can compensate for missing motion cues with finer frame details.
Second, removing segmentation (w/o Seg, row 1) hurts high-bitrate (Level 3) LPIPS most compared with the full model (row 1 vs. row 4), as segmentation preserves contours (e.g., text/rigid objects) that become perceptually critical when other artifacts are minimized. At low bitrates (Level 1), its impact is weaker because coarse compression overwhelms segmentation-guided details. The benefits of optical flow (w/o Flow, row 3) are robust across different bitrate levels, showing consistent gains for dynamic content, especially in complex motion (e.g., crowd scenes).
Last, the complete model (all conditions, row 4) achieves the best results, outperforming all other settings. This confirms that conditions are complementary: Motion anchors high-level dynamics, while Segmentation and Flow refine spatial and local motion details.
Visualized Results
Figure 5: Visual comparison with state-of-the-art codec on the test set. For each case, we present the frames (with an interval of 7 frames) between original video (top row), our model (middle row) and H.264 codec (bottom row) . Both H.264 compression and the videos compressed by our model are controlled to have a bpp of 0.0066 (Best viewed zoomed in and in color).
In this subsection, we present some visualized results to demonstrate the performance of our model and comparative methods intuitively.
Fig. 5 provides a qualitative comparison of video compression performance across different methods, specifically between Ours, H.264, and the Ground Truth. The selected samples illustrate how our method preserves semantic and structural integrity under four complex scenarios.
(a) Scene: A Young Man on a Cliff
In this natural outdoor scene, the Ground Truth image contains rich texture details, including the rocks, the ocean, and the subject’s facial features. Our method closely resembles the Ground Truth, successfully preserving sharp edges and natural colors in the background. H.264, on the other hand, introduces visible artifacts and blurring, particularly around the edges of the subject and the textures of the rocks. The superior detail retention of our approach highlights its capability in handling complex natural scenes with high-frequency textures.
(b) Scene: Person Fishing on Water
This case involves dynamic content, including water reflections and motion. The Ground Truth shows clear water ripples and fine details in the fishing gear and clothing textures. Our method again demonstrates superior visual quality, preserving both the clothing structure and the rippling water with minimal degradation. In contrast, H.264 results in a smeared appearance, especially in the water region, where fine-grained textures are lost. This demonstrates our model’s robustness in maintaining fidelity in motion-heavy, fine-detailed scenes.
(c) Scene: Indoor Interview
This indoor scenario introduces artificial lighting and facial features in a human-centric conversation setup. The Ground Truth image preserves skin tone consistency and clothing folds. Our method produces sharp and visually appealing results that retain fine facial details and contrast. H.264 fails to reproduce the same level of clarity, with noticeable blur in facial features and clothing edges leading to a loss of expression and scene semantics. The results emphasize our model’s advantage in retaining critical human visual cues at ultra low bitrates, even under indoor lighting.
(d) Scene: TEDx Presentation
This example includes stage lighting, text (TEDx letters), and human presence. The Ground Truth shows clear red letters and consistent skin tone under spotlight. Our method maintains high fidelity in both the speaker’s outline and the textured stage background, attribute to segmentation sequences and human motion. The H.264 result suffers from color bleeding and edge artifacts, especially around the large red letters and the person’s silhouette. This shows our model’s capacity to preserve both foreground and background details with high structural accuracy, even under challenging lighting conditions.
In summary, across all four scenes, our method consistently outperforms H.264 by delivering higher perceptual quality and better semantic integrity at ultra low bitrates, which confirms that our method can provide high-quality compressed outputs suitable for applications with heavy transmission constraints, such as satellite video conferencing, surveillance, and media streaming.
Conclusion
This paper presents the first end-to-end video compression framework leveraging conditional diffusion models to achieve human-perception-aligned reconstruction at ultra-low bitrates. Despite these advances, limitations remain: The current decoding speed falls short of real-time requirements due to computational intensity in diffusion-based generation. To address this, we propose adopting familiar model—deploying homologous diffusion models of varying capacities—enabling automatic model switching based on decoder-side computational resources, inspired by AI Flow paradigms (An et al., 2025). Additionally, content-agnostic bitrate control via manual parameter tuning (e.g., $l$, NN, $\xi$) limits adaptability. Future work will integrate content-understanding modules to dynamically optimize compression thresholds based on scene complexity and motion characteristics.